FEL Upgrade Project Weekly Brief
July 31 Aug. 4, 2000


This week during IR Demo operations we characterized the angular and energy distribution of the short pulse x-rays produced by Compton scattering and delivered light for experiments on protein dynamics, carbon nanotube
production and polymer ablation.  For the upgrade, the order for the high power circulators for the upgraded injector has been placed.


This is the last week of the summer run for FEL users.  Next week attention turns to summarizing our upgrade design activities and the FEL operation results for the International FEL Conference which will be held at Duke University (Aug 14-18) and for the Linac 2000 Conference which will be held the following week (Aug.21-25) at Monterey.

We sent an invitation to RADM M. Mathis, Commander NAVSEA, per a request from his office for a tour and briefing on the JLab FEL Program.  On Friday, Aug. 4, JLab hosted a contingent from OMB, led by Gary Bennethum, who
oversees many of the DOE science and technology programs.

WBS 3 (Beam Physics)

WBS 3 activities centered on preparation for FEL 2000 and Linac 2000 and on planning design and beam dynamics studies over the duration of the project.  A list of machine physics issues and timeline for their resolution was
established, and interest in participation in studies on particular topics declared.

WBS 4 (Injector)

We are trying to rid a heated test chamber and cesiator of hydrocarbons using a hot nitrogen purge under moderate vacuum maintained by a turbopump.  If this works, we will be able to address if the source of our anomalous
high pressure during bake out.


Horizontal Test Bed, HTB, activities have included cold testing of mechanical and piezo tuner system.  Tuner performance meets all specifications. Microphonic and rf coupling testing continues this week. Second generation, cost optimized, tuner procurement has been placed. Helium vessel procurement placed, vendor visit scheduled for 5 September. Vacuum vessel and space frame procurement being finalized, vendor selection complete.  Procurement will be finalized next Monday and a vendor visit will be scheduled at that time.  Waveguide purchase requests completed.  End can SOW and drawing package complete, purchase request will be finalized next week.

WBS 6 (RF)

We have placed the order for the high power (100 kW) circulators that we need for the upgrade injector RF power systems.  Three units (2 plus a spare) have been ordered in a competitive procurement won by AFT, Inc.

WBS 8  (I & C)

The following orders were placed:  1) IR laser diode and driver from Thor Labs for continued development of on line optical diagnostics.  2) Two 32 element linear pyroelectric arrays were received - one with and one without
a window.  3) Quotes are coming in for the ion pump power supply components.  First cut at the numbers look like 25 for IR upgrade and 75 for SNS; this will greatly improve the prices we will pay.  4) Discussions have been started with other labs about trim power supplies.  The cost and reliability (including spares & CAMAC) of our trim racks have prompted a
good look at alternatives.  5) Fabricated and installed an X-Y translator stage to be able to track the X-Ray detector with the crystal, this worked very well.  6) A 60Hz trigger distribution chassis was completed (4 X 1:4) to distribute the drive laser (& FEL) relative phase reference to multiple locked optical choppers on the FEL diagnostics beam line.

WBS 9 (Transport)

Beam Transport System (WBS 9)
o  We continued layout of the GX Prototype further refining the cooling water the internal bus bar connections, external       lead connections coil clamping and cooling hose configuration.
o  The completed 3D magnetic model of the GX dipole yielded results showing that the good field region is within specification at the mid plane.  Other specifications await further analysis of the data.

o  Start of finish machining of one steel quadrant with CNC is delayed to Monday, August 14th because the tooling required stiffening.
o  The first coil of the prototype quadrupole wound.

o  Proposals for the engineering services contract are near the end of technical evaluation.
o  Overall layouts of the machine configuration continue.  The first of the drawings, one that defines the regions of the machine is ready for signoff.
o  We feel confident that we can eliminate a potential interference between the flanges of the larger diameter Beam Position Monitor and the dipole correctors, which we wanted to reuse from the IR Demo.

WBS 10 (Wiggler)

Coils have been installed on the first core of the Northrop Grumman Wiggler.  The vacuum pipe drawing is ready for checking while design of the diagnostics and water manifolds continues.

Drawings for the dispersion section were modified to ensure good gap uniformity and are ready for sign-off.

WBS 11 (Optics)

Most of our efforts this week were devoted to operations issues mentioned elsewhere.  We did initiate discussions with other optical scientists to learn specifics on the low-loss coatings developed for the ABL and SBL programs. Some discussions are also taking place with TRW staff with the status of their modeling of the R5 resonator design. Incremental discussions continue with vendors on additional optical diagnostics, and we have received a sensitive pyroelectric array for evaluation.


This was the last week of the summer user run.  As noted in last week's report FEL operations were plagued with an intermittent phase stability problem somewhere between the master oscillator (MO) and the drive laser mode locker.  The problem first started to affect operations last Monday (July 24th) and we didn't eliminate the problem until this past Tuesday.  We spent most of last Saturday through Tuesday am diagnosing the problem with a top-to-bottom look at all the RF hardware connected to the drive laser.  (We thank Curt Hovater and John Mussen for their support on this problem).  The problem was traced to several RF connections that had gone sour in the circuitry which divides the MO signal by 40 prior to the laser interface.  Since the problem was repaired on Tuesday the FEL has run with high stability and reliability.

The principal user this week was Bob Austin from Princeton University who is continuing his studies of protein dynamics at the amide I band at 6 microns.  As of this writing Bob has spent 6 shifts checking out a new experimental set-up and has not yet collected any new pump-probe spectra because of an undiagnosed problem in at least one optical element in his

We have been able to piggy-back three other experiments this week.  The first runs with a new user set-up from a CWM-PSU-NASA team was tested to produce carbon nanotubes by carbon laser ablation.  On the second attempt,
significant soot like product was made with 400 watts of FEL power.  The soot has been forwarded to PSU for Raman/TEM analysis to confirm the presence of nanotubes.  An attempt was made to continue M. Kelley's
experiments on polymer ablation at 6 microns but the initial micropulse energy was too small (< 3.5 microjoules) and not enough time was available to optimize at higher pulse energies.

By piggybacking during last weekend's diagnostic runs and during night owl shifts this week, some very nice results were obtained with the Compton x-ray experiment.

We started characterizing the spatial distribution of the x-rays.  An important initial result is that imaging cameras - based on position sensitive scintillators - cannot be used for this task.  There are too many background x-rays (of all energies) that light up the scintillator.  Energy discrimination must be used.  Consequently, the Si PIN diode used to
initially detect x-rays last Sept. was mounted on a x-y translation stage and we successfully began mapping the distribution. On the plus side, at the energy of the Thomson x-rays, the signal to noise ratio is about 90-to-1.  One puzzler developed: the distribution seems to be about a factor of four sharper than the 1/gamma prediction.  These initial results will be shared
with collaborators at the upcoming FEL2000 conference at Duke University in a couple of weeks.

Privacy and Security Notice